Assume the probability it will rain today is 0.634. Calculate the probability it will not rain today.

A jar of marbles contains 4 red marbles, 8 blue marbles and 10 black marbles.

a. What is the probability I randomly select one marble and it is blue?

(4 points)

b. What is the probability I randomly select one marble and it is not red?

(4points)

c. What is the probability I randomly select on marble and it is blue or black?

(6 points)
Answer #1

2 Total ro of Marbles z (4+ 8 +10) Z 22 P (Randomly selected Marble ix blue) NO OF Blue Moubles Total No of Masubles = 8 = 0.P(Blue or block) - P (Blue) + P (block) - P(Blue and block) + 10 - p (awe)xp (bloek) 22 = 0.3636 + 0.4545 – 0.1653, = 0.6528Note-if there is any understanding problem regarding this please feel free to ask via comment box..thank you